How do you write and solve a two step equation? i am in pre algebra NOT ALGEBRA!
for example:
eilleen swam for 85 minutes. this is 21 more minutes than 4 times the number of minutes Ethan swam. i thought the answer might be
85-21 divided by 4
thank you for your help!
IF YOU COULD E-MAIL ME AS SOON AS PODSSIBLE BECAUSE THIS IS MY HOMEWORK AND I HAVE TO TURN IT IN TOMARROW!!!

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

eilleen swam for 85 minutes. this is 21 more minutes than 4 times the number of minutes Ethan swam. i thought the answer might be
85-21 divided by 4
----------------------
That's what it is.
(85-21)/4 = 64/4 = 16 minutes
--------------------
To check: 16*4 + 21 = 85

Answer by richwmiller(17219)

You are correct.
You didn't state the whole problem.
I will assume you want to know how long Ethan swam. There are other possible questions.
N=time Ethan swam
L=85
85=21+4n
and your solution comes in next
Subtract 21 from both sides
85-21=4n
simplify
64=4n
divide both sides by 4
16=n
Ethan swam 16 minutes.